OpenStudy (jazzyfa30):

The football team has 30 more members than the basketball teams. If each group had 10 more members, the ratio of their membership would be 3:2. How many members are in each group? Write your answer as football x basketball y

OpenStudy (anonymous):

x=y+30 2(x+10)=3(y+10) hence 2x+20=3y+30 2(y+30)+20=3y+30 2y+80=3y+30 y=50 x=80

OpenStudy (jazzyfa30):

On a particular day, the wind added 3 miles per hour to Alfonso's rate when he was cycling with the wind and subtracted 3 miles per hour from his rate on his return trip. Alfonso found that in the same amount of time he could cycle 36 miles with the wind, he could go only 24 miles against the wind.What is his normal bicycling speed with no wind? (Remember, speed = distance/time) @bahrom7893

OpenStudy (jazzyfa30):

@mathstudent55

OpenStudy (anonymous):

put yo boobiehs away girl!

OpenStudy (jazzyfa30):

lol thanks for noticing but can u help me

OpenStudy (anonymous):

Im thinking

OpenStudy (jazzyfa30):

ummmmmmmmmm i guess u dont know

OpenStudy (anonymous):

nope im stuck

OpenStudy (jazzyfa30):

r u sure ur not distracted by my breast

OpenStudy (anonymous):

im sure lol hehehe

OpenStudy (jazzyfa30):

lol u say that

OpenStudy (jazzyfa30):

@matimatics

OpenStudy (jazzyfa30):

@Jhannybean

OpenStudy (anonymous):

time1 = time2 24 miles / (v - 3 mph) = 36 miles / (v + 3 mph) 24 * (v+3) = 36 * (v-3) 24v + 72 = 36v - 108 12v = 180 v=15 mph

OpenStudy (jazzyfa30):

\[\frac{ 1 }{ x-1 }+ \frac{ 2 }{ x }=0\]

OpenStudy (anonymous):